The sum of the first and third of three consecutive even integers is 176. What are the three even integers?

Answer by stanbon(75887)

The sum of the first and third of three consecutive even integers is 176. What are the three even integers?
---
1st: 2x-2
2nd: 2x
3rd: 2x+2
-----
Equation:
2x-2 + 2x+2 = 4x
---
4x = 176
x = 44
-----
1st: 2x-2 = 86
2nd: 2x = 88
e4d: 2x+2 = 90
----------
Cheers,
Stan H.